Logistics Solutions provides order fulfillment services for dot.com merchants. The company maintains warehouses that stock items carried by its dot.com clients. When a client receives an order from a customer, the order is forwarded to Logistics Solutions, which pulls the item from storage, packs it, and ships it to the customer. The company uses a predetermined variable overhead rate based on direct labor-hours. In the most recent month, 190,000 items were shipped to customers using 8,300 direct labor-hours. The company incurred a total of $29,050 in variable overhead costs. According to the company’s standards, 0.03 direct labor-hours are required to fulfill an order for one item and the variable overhead rate is $3.55 per direct labor-hour. Required:

1. What is the standard labor-hours allowed (SH) to ship 190,000 items to customers?

2. What is the standard variable overhead cost allowed (SH × SR) to ship 190,000 items to customers?

3. What is the variable overhead spending variance?

4. What is the variable overhead rate variance and the variable overhead efficiency variance?
Answer #1

1. Standard labor-hours allowed (SH) = 190000 x 0.03 = 5700

2. Standard variable overhead cost allowed (SH x SR) = 5700 x $3.55 = $20235

3. Variable overhead spending variance = Actual overhead costs - Standard overhead cost = $29050 - $20235 = $8815 Unfavorable

4. Variable overhead rate variance = (AH x AR) - (AH x SR) = $29050 - (8300 x $3.55) = $29050 - $29465 = $415 Favorable

Variable overhead efficiency variance = SR x (AH - SH) = $3.55 x (8300 - 5700) = $3.55 x 2600 = $9230 Unfavorable